Construct a table of valued of function using the interval of -5 to 5.

j(t)=1/(t²-2t+1)


1
Expert's answer
2021-09-05T18:15:46-0400
"j(t)=\dfrac{1}{t^2-2t+1}"

"j(t)=\dfrac{1}{(t-1)^2}"

"t\not=1"

Doman: "(-\infin, 1)\cup(1, \infin)."

If we consider the interval "[-5, 5]," then "t\in[-5, 1)\cup(1, 5]."


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 t & y=j(t) \\ \hline
 -5 & 1/36 \\
 \hdashline
 -4.5 & 4/121 \\
 \hdashline
 -4 & 1/25 \\
 \hdashline
 -3.5 & 4/81 \\
 \hdashline
 -3 & 1/16 \\
 \hdashline
 -2.5 & 4/49 \\
 \hdashline
 -2 & 1/9 \\
 \hdashline
 -1.5 & 4/25 \\
 \hdashline
 -1 & 1/4 \\
 \hdashline
 -0.5 & 4/9 \\
 \hdashline
0 & 1 \\
 \hdashline
 0.5 & 4 \\
 \hdashline
 0.6 & 25/4 \\
 \hdashline
 0.7 & 100/9 \\
 \hdashline
 0.8 & 25 \\
 \hdashline
 0.9 & 100 \\
 \hdashline
 1.1 & 100 \\
 \hdashline
 1.2 & 25\\
 \hdashline
 1.3 & 100/9 \\
 \hdashline
 1.4 & 25/4 \\
 \hdashline
 1.5 & 4 \\
 \hdashline
 2 & 1 \\
 \hdashline
2.5 & 4/9 \\
 \hdashline
3 & 1/4 \\
 \hdashline
 3.5 & 4/25 \\
 \hdashline
 4 & 1/9 \\
 \hdashline
4.5 & 4/49 \\
 \hdashline
 5 & 1/16 \\

\end{array}"